OpenStudy (anonymous):

Find the area of the biggest possible square that would fit into a circle having a radius of 3cm.

OpenStudy (anonymous):

I would say 16 cm but I am not sure, as I do not entirely know if the space between the side of the square and the farthest point on the circle from it is 1 cm. See if someone else will confirm if I am right or wrong.

OpenStudy (mathmate):

 Using Pythagoras theorem, \(x^2=3^2+3^2=area~of~square\)